25x+5y=38
-5x=y+10
solve by substitution method

Answer by stanbon(75887)

25x+5y=38
-5x=y+10
solve by substitution method
------------------------------------
y = -5x-10
--------------
Substitute for "y" and solve for "x":
25x + 5(-5x-10) = 38
25x -25x - 50 = 38
-50 = 38
----------
That is a contradiction
The equation is inconsistent.
There are no solutions.
Cheers,
Stan H.